# Double rotatory power reversal and continuous Kerr angle in bi-isotropic media with anomalous Hall current

### v1

@@ Abstract
We investigate optical properties of bi-isotropic materials under the anomalous Hall effect (AHE) of the axion electrodynamics. Four refractive indices associated with circularly polarized waves are achieved, implying circular birefringence with rotatory power (RP) endowed with double sign reversal, an exotic optical signature for chiral dielectrics. The Kerr rotation and ellipticity are analyzed, with an unusual observation of rotation angle deprived of discontinuity. Anomalous reflection (greater than unity) is also reported, associated with negative refraction stemming from the anomalous transport properties. These effects constitute the singular optical signature of bi-isotropic media with the AHE.
@@
Introduction. Optical characterization of chiral media constitutes a powerful technique for matter characterization. Birefringence appears as a key feature of optically active media, being measured by the rotatory power (RP) [1], which finds application in the description of bi-isotropic matter [2, 3], organic compounds [4], and chiral metamaterials [5]. The dispersive RP can be anomalous if it undergoes reversion with the frequency [6]. The Kerr rotation angles provide magneto-optical signatures of reflected light [7, 8]. The magneto-optic Kerr effect (MOKE) and magneto-optical effects in complex materials are broadly employed to probe properties of topological insulators [9], new graphene composites [10] and to examine giant Kerr rotation in Weyl semimetals [11] and time-reversal broken systems [12].
@@
Nonconventional phenomena, such as the Chiral Magnetic Effect (CME), where an electric current is generated along an applied magnetic field due to an imbalance in the number density of chiral fermions with opposite handedness [13, 14], has played an important role in Weyl metals [15] and semimetals [16, 17]. Another relevant quantum-anomaly induced transport phenomenon is the anomalous Hall effect (AHE), due to the separation of Weyl nodes in momentum for right- and left-handed fermions. The AHE and the CME can be effectively described by including the axion term [20, 21, 18, 19], $\mathcal{L}=\theta(\mathbf{E}\cdot\mathbf{B})$, in the classical Maxwell Lagrangian in continuous matter, where the anomalous Hall current takes place, $\nabla\theta\times\mathbf{E}$. For the special case of a nondynamical axion field, $\partial_{t}\theta=cte$, $\nabla\theta=cte$, one recovers the Maxwell-Carroll-Field-Jackiw (MCFJ) electrodynamics [22] in continuous matter [23, 24]. For a time-independent axion field, $\partial_{t}\theta=0$, the Ampere’s law reads ×$\nabla\times\mathbf{H}-\partial_{t}\mathbf{D}=\mathbf{J}+\mathbf{b}\times\mathbf{E}$, with $\mathbf{b}=\nabla\theta$ standing for the axion field gradient. The axion electrodynamics effectively describes relevant aspects of Weyl semimetals [25], optical properties of exotic metamaterials [26], axion dielectrics [27, 28], connections with the London equation and Weyl semimetals [29, 30], and optical reflection properties at the surface of an axion dielectric [31]. Wave propagation in bi-isotropic media endowed with the axion magnetic current term was classically addressed as well, yielding dispersive birefringence and rotatory power endowed with sign reversal [32], a typical property of rotating plasmas [33], chiral plasmas [34], graphene systems [35], and particular Weyl semimetals [36]. Investigations on reflection for normal incidence at a Weyl semimetal surface have yielded anomalous reflectance (greater than unity) due to chiral magnetic instabilities associated with CME and AHE axion terms [37, 38]. Moreover, nonreciprocal thermal radiation emission in Weyl semimetals (with AHE) has shown negative emissivity at low frequencies, deviating from Planck’s law [39]. Kerr rotation and ellipticity angles [40] depending on the frequency were reported in Weyl semimetals [41], where it can be used to design circular polarizers or optical isolators [42, 43].
@@
In this letter, we investigate the optical characterization of bi-isotropic media endowed with the AHE, including circular birefringence, Kerr rotation, Kerr ellipticity, and anomalous reflection, achieving new remarkable features that compose a specific optical signature of such a system. Throughout this work, we use natural units.
@@
Dispersion relations and refractive indices. In the plane wave ansatz, one writes the Maxwell equations modified by the AHE term,
$\mathbf{k}\cdot\mathbf{D}=-\mathbf{b}\cdot\mathbf{B},\quad\quad\mathbf{k}\times\mathbf{H}=-\omega\mathbf{D}+i\,\mathbf{b}\times\mathbf{E},$ (1)
@@
$n_{RCP,\pm}$ $=\mu\alpha^{\prime\prime}\pm Q_{-},\quad n_{LCP,\pm}=-\mu\alpha^{\prime\prime}\pm Q_{+},$ (4)
$Q_{\pm}$ $=\sqrt{\mu\epsilon+\mu^{2}\alpha^{\prime\prime 2}\pm\mu b/\omega}.$ (5)
where $\mathbf{n}$ is considered parallel to the vector $\mathbf{b}$, and $\alpha+\beta=2i\alpha^{\prime\prime}$, with $\mathrm{Im}[\alpha]=\alpha^{\prime\prime}$. The indices $n_{RCP,\pm}$ and $n_{LCP,\pm}$ are associated with right-handed circularly polarized (RCP) mode and left-handed circularly polarized (LCP) wave, respectively, given by $\hat{\bf{e}}_{RCP,\,LCP}=\hat{\bf{x}}\mp i\hat{\bf{y}}$.
The indices $n_{RCP,\pm}$ are complex in the frequency window $\omega<\omega_{0}$, with
@@
being $0<\omega<\omega_{0}$ the absorption zone (for positive definite electromagnetic parameters). In Fig. 1, we illustrate the frequency behavior of the indices $n_{RCP,\pm}$ of Eq. (4). For $\omega>\omega_{0}$, the indices become real, being $n_{RCP,+}$ (red line) positive and monotonically increasing (normal dispersion), while $n_{RCP,-}$ (blue line) is progressively decreasing (anomalous dispersion).
Figure 1: Refractive indices $n_{RCP,\pm}$ of Eq. (4). The red (blue) lines represent $n_{RCP,+}$ ($n_{RCP,-}$). The solid (dashed) line indicates the real (imaginary) part of the refractive indices. The solid purple line represents the real parts of $n_{RCP,+}$ and $n_{RCP,-}$, lying on top of each other. Here, we have used $\mu=1,\epsilon=3,\alpha^{\prime\prime}=3$ and $b=1$ $\mathrm{s}^{-1}$.
Figure 2: Refractive indices $n_{LCP,\pm}$ of Eq. (4). The red (blue) line indicates the positive (negative) refractive index ($n_{LCP,\pm}$), respectively. The horizontal dashed lines are given by Eq. (7). Here, we have used $\mu=1,\epsilon=3,\alpha^{\prime\prime}=3$ and $b=1$ $\mathrm{s}^{-1}$.
Furthermore, the index $n_{RCP,-}$ becomes negative for $\omega>\omega_{R}$ with $\omega_{R}=b/\epsilon$, as shown in Fig. 1. As for the LCP modes, the associated refractive indices, $n_{LCP,\pm}$, are both real, being $n_{LCP,-}$ negative for any frequency, while $n_{LCP,+}$ is always positive, as shown in Fig. 2. In the high-frequency regime, the indices $n_{LCP,\pm}$ and $n_{RCP,\pm}$ tend to positive and negative asymptotic values,
$\left.n_{LCP,\pm}\right|_{\mathrm{high}}$ $=-\mu\alpha^{\prime\prime}\pm\sqrt{\mu\epsilon+\mu^{2}\alpha^{\prime\prime 2}},$ (7)
$\left.n_{RCP,\pm}\right|_{\mathrm{high}}$ $=\mu\alpha^{\prime\prime}\pm\sqrt{\mu\epsilon+\mu^{2}\alpha^{\prime\prime 2}},$ (8)
@@
Birefringence and Rotatory Power. The circular birefringence is measured in terms of the rotatory power,
$\delta_{RP}=-\dfrac{\omega}{2}\left[\operatorname{Re}(n_{LCP,\pm})\ -\operatorname{Re}(n_{RCP,\pm})\right].$ (9)
The fact that we have two refractive indices for each propagating wave indicates that we have the possibility of four distinct results for the rotatory power, given as follows:
@@
$\delta_{-,\pm}$ $=\frac{\omega}{2}\operatorname{Re}\left(2\mu\alpha^{\prime\prime}+Q_{+}\pm Q_{-}\right),$ (11)
One notes that $Q_{-}$ contributes only when it is real, that is, for $\omega>\omega_{0}$, with $\omega_{0}$ given by Eq. (6). The behavior of the rotatory power $\delta_{+,\pm}$ of Eq. (10) as a function of the frequency can be observed in Fig. 3, being both equal in the range $0<\omega<\omega_{0}$ (see the purple line), where they receive contribution only of the first root ($Q_{+}$). In this frequency range, both RPs $\delta_{+,\pm}$ possess sign reversal at $\omega=\omega^{\prime}$, with $\omega^{\prime}$ given by $\omega^{\prime}={b}/({3\mu\alpha^{\prime\prime 2}-\epsilon})$.
@@
For $\omega>\omega_{0}$, the RPs begin to receive the contribution of the second root, resulting in the splitting starting at $\omega=\omega_{0}$, from which $\delta_{+,+}$ exhibits an increasing behavior (see the red line), while $\delta_{+,-}$ has decreasing magnitude (see the blue line). Along with that, the RP $\delta_{+,-}$ exhibits a second (additional) sign reversal at the frequency, $\omega^{\prime\prime}={b}/({2\alpha^{\prime\prime}\sqrt{\mu\epsilon}})$, with $\omega^{\prime\prime}>\omega_{0}$. This double sign inversion is an effect of both the magnetoelectric and AHE parameters, a very unusual feature that may work as an optical signature of this chiral system. A single RP reversion was reported for the case of a bi-isotropic dielectric in the presence of an isotropic axion magnetic current [32]. For high-frequencies, where the term $b/\omega$ becomes negligible, the coefficients $\delta_{+,+}$ and $\delta_{+,-}$ recover a linear behavior with a positive (negative) slope.
@@
Figure 3: Rotatory power $\delta_{+,\pm}$ of Eq. (10), constructed using the two refractive indices for the RCP wave and $n_{LCP,+}$. The RP $\delta_{+,+}$ is depicted by the purple-red line, while $\delta_{+,-}$ is demarcated by the purple-blue line, endowed with double sign reversal. Here, we have used $\mu=1,\epsilon=\alpha^{\prime\prime}=3$, and $b=1$ $\mathrm{s}^{-1}$.
@@
Kerr rotation and Kerr ellipticity. Considering incident light stemming from medium 1 (a usual dielectric described by permittivity $\epsilon_{1}$ and permeability $\mu_{1}$) and reflecting on the surface of medium 2 (a chiral dielectric described by the parameters $\epsilon$, $\mu$, the axion vector $\mathbf{b}$, and the refractive indices of Eq. (4)), we obtain two Fresnel coefficients for each left-handed circularly polarized (LCP) mode and right-handed circularly polarized (RCP) wave at normal incidence, namely
$r_{LCP,\pm}=\frac{\mu_{1}n_{LCP,\pm}-\mu n_{1}}{\mu_{1}n_{LCP,\pm}+\mu n_{1}},$ (12a)
$r_{RCP,\pm}=\frac{\mu_{1}n_{RCP,\pm}-\mu n_{1}}{\mu_{1}n_{RCP,\pm}+\mu n_{1}}.$ (12b)
The latter can be used to define the Kerr angle expression
$\Delta_{\pm,\pm}$ $=\left(\frac{r_{LCP,\pm}-r_{RCP,\pm}}{r_{LCP,\pm}+r_{RCP,\pm}}\right),$ (13)
@@
where the first (second) $\pm$ subscript is related to $n_{LCP,\pm}$ ($n_{RCP,\pm}$). The elements $\Delta_{\pm,\pm}$ are generally complex,
$\Delta_{\pm,\pm}=\Delta_{\pm,\pm}^{\prime}+i\Delta_{\pm,\pm}^{\prime\prime},$ (14)

### v2

# Double rotatory power reversal, continuous Kerr angle, and enhanced reflectance in bi-isotropic media with anomalous Hall current

@@ Abstract
We investigate the optical properties of bi-isotropic materials under the anomalous Hall effect (AHE) of axion electrodynamics. Four refractive indices associated with circularly polarized waves are achieved, implying circular birefringence with rotatory power endowed with double sign reversal, an exotic optical signature for chiral dielectrics. The Kerr rotation and ellipticity are analyzed, with an unusual observation of a giant rotation angle deprived of discontinuity. Anomalous enhanced reflectance (R greater than unity) is also reported, associated with negative refraction stemming from the anomalous transport properties. These effects constitute the singular optical signature of a nonequilibrium bi-isotropic medium with the AHE.
I Introduction
@@
The optical characterization of chiral media constitutes a powerful technique for matter characterization. Birefringence appears as a key feature of optically active media, being measured by the rotatory power (RP) [1], which finds application in a broad set of scenarios, including bi-isotropic matter [2, 3], organic compounds [4], chiral metamaterials [5], graphene based devices [6], sensors and filters [7], chiral metasurfaces [8], polarization rotators [9], dielectric and quasi-planar nanostructures [10, 11]. The dispersive RP can be anomalous if it undergoes reversion with the frequency [12]. The Kerr rotation angles provide magneto-optical signatures of reflected light [13, 14]. The magneto-optic Kerr effect (MOKE) and magneto-optical effects in complex materials are broadly employed to probe properties of topological insulators [15], new graphene composites [16] and to examine giant Kerr rotation in Weyl semimetals [17] and time-reversal broken systems [18].
@@
Nonconventional phenomena, such as the Chiral Magnetic Effect (CME), where an electric current is generated along an applied magnetic field due to an imbalance in the number density of chiral fermions with opposite handedness [19, 20], has played an important role in Weyl metals [21] and semimetals [22, 23]. Another relevant quantum-anomaly induced transport phenomenon is the anomalous Hall effect (AHE), due to the separation of Weyl nodes in momentum for right- and left-handed fermions. The AHE and the CME can be effectively described by including the axion term [26, 27, 24, 25], $\mathcal{L}=\theta(\mathbf{E}\cdot\mathbf{B})$, in the classical Maxwell Lagrangian in continuous matter, where the anomalous Hall current takes place, $\nabla\theta\times\mathbf{E}$. For the special case of a nondynamical axion field, $\partial_{t}\theta=cte$, $\nabla\theta=cte$, one recovers the Maxwell-Carroll-Field-Jackiw (MCFJ) electrodynamics [28] in continuous matter [29, 30]. For a time-independent axion field, $\partial_{t}\theta=0$, the Ampere’s law reads $\nabla\times\mathbf{H}-\partial_{t}\mathbf{D}=\mathbf{J}+\mathbf{b}\times\mathbf{E}$, with $\mathbf{b}=\nabla\theta$ standing for the axion field gradient. The axion electrodynamics effectively describes relevant aspects of Weyl semimetals [31, 32], optical properties of exotic metamaterials [33], axion dielectrics [34, 35, 36], connections with the London equation and Weyl semimetals [37, 38], Cherenkov radiation [39], applications in ultrafast magnetism [40], photonics of new chiral materials [41], and optical reflection properties at the surface of an axion dielectric [42]. Wave propagation in bi-isotropic media endowed with the axion magnetic current term was classically addressed as well, yielding dispersive birefringence and rotatory power endowed with sign reversal [43], a typical property of rotating plasmas [44], chiral plasmas [45], graphene systems [46], and particular Weyl semimetals [47]. Investigations on reflection for normal incidence at a Weyl semimetal surface have yielded anomalous reflectance (greater than unity) due to chiral magnetic instabilities associated with CME and AHE axion terms [48, 49, 50]. Moreover, nonreciprocal thermal radiation emission in Weyl semimetals (with AHE) has shown negative emissivity at low frequencies, deviating from Planck’s law [51]. Kerr rotation and ellipticity angles [52] depending on the frequency were reported in Weyl semimetals [53], where it can be used to design circular polarizers or optical isolators [54, 55].
@@
In this paper, we investigate the optical characterization of bi-isotropic media endowed with the AHE, including circular birefringence, Kerr rotation, Kerr ellipticity, and anomalous reflection, achieving new remarkable features that compose a specific optical signature of such a system. Throughout this work, we use natural units.
II Dispersion relations and refractive indices
@@
In the plane wave ansatz, one writes the Maxwell equations modified by the AHE term,
$\mathbf{k}\cdot\mathbf{D}=-i\mathbf{b}\cdot\mathbf{B},\quad\quad\mathbf{k}\times\mathbf{H}=-\omega\mathbf{D}+i\,\mathbf{b}\times\mathbf{E},$ (1)
@@
$n_{1,\pm}$ $=\mu\alpha^{\prime\prime}\pm Q_{-},\quad n_{2,\pm}=-\mu\alpha^{\prime\prime}\pm Q_{+},$ (4)
where $\mathbf{n}$ is considered parallel to the vector $\mathbf{b}$,
$Q_{\pm}=\sqrt{\mu\epsilon+\mu^{2}\alpha^{\prime\prime 2}\pm\mu b/\omega},$ (5)
and $\alpha+\beta=2i\alpha^{\prime\prime}$, with $\mathrm{Im}[\alpha]=\alpha^{\prime\prime}$. The propagation modes for waves along the z-axis are described by the polarization vectors $\hat{\bf{e}}_{\mp}=\hat{\bf{x}}\mp i\hat{\bf{y}}$ – right-handed circularly polarized (RCP) and left-handed circularly polarized (LCP) waves, respectively. The two indices $n_{1,\pm}$ are associated with the RCP polarization vector ($\hat{\bf{e}}_{-}$). While $n_{1,+}$ designates forward RCP propagation for any frequency, the negative index $n_{1,-}$ corresponds to forward RCP propagation in the frequency range $\omega<\omega_{R}$ and backward LCP propagation for $\omega>\omega_{R}$, where $\omega_{R}=b/\epsilon$. The indices $n_{2,\pm}$ are associated with the LCP vector ($\hat{\bf{e}}_{+}$), with $n_{2,+}$ indicating forward LCP propagation and $n_{2,-}$ a backward RCP wave. The four solutions in Eq. (4) are distinct from one another. Along with that, one observes that the propagation of RCP forward and backward waves are associated with two distinct refractive indices (when considering a global minus sign, due to $\pm\mu b/\omega$ factor in $Q_{\pm}$). The same is true for the forward and backward LCP waves. This is a manifestation of nonreciprocity (stemming from the AHE term).
The indices $n_{1,\pm}$ are complex in the frequency range $\omega<\omega_{0}$, with
@@
where $0<\omega<\omega_{0}$ is the absorption zone (for positive definite electromagnetic parameters). In Fig. 1, we illustrate the frequency behavior of the indices $n_{1,\pm}$ of Eq. (4). For $\omega>\omega_{0}$, the indices become real, being $n_{1,+}$ (red line) positive and monotonically increasing (normal dispersion), while $n_{1,-}$ (blue line) is progressively decreasing (anomalous dispersion).
Figure 1: Refractive indices $n_{1,\pm}$ of Eq. (4). The red (blue) lines represent $n_{1,+}$ ($n_{1,-}$). The solid (dashed) line indicates the real (imaginary) part of the refractive indices. The solid purple line represents the real parts of $n_{1,+}$ and $n_{1,-}$, lying on top of each other. Here, we have used $\mu=1,\epsilon=3,\alpha^{\prime\prime}=3$ and $b=1$ $\mathrm{s}^{-1}$.
Figure 2: Refractive indices $n_{2,\pm}$ of Eq. (4). The red (blue) line indicates the positive (negative) refractive index ($n_{2,\pm}$), respectively. The horizontal dashed lines are given by Eq. (8). Here, we have used $\mu=1,\epsilon=3,\alpha^{\prime\prime}=3$ and $b=1$ $\mathrm{s}^{-1}$.
Furthermore, the index $n_{1,-}$ becomes negative for $\omega>\omega_{R}$, with $\omega_{R}=b/\epsilon$, as shown in Fig. 1. Also, the refractive indices $n_{2,\pm}$ are both real, being $n_{2,-}$ negative for any frequency, while $n_{2,+}$ is always positive, as shown in Fig. 2. In the high-frequency regime, the indices $n_{1,\pm}$ and $n_{2,\pm}$ tend to positive and negative asymptotic values,
$\left.n_{1,\pm}\right|_{\mathrm{high}}$ $=\mu\alpha^{\prime\prime}\pm\sqrt{\mu\epsilon+\mu^{2}\alpha^{\prime\prime 2}},$ (7)
$\left.n_{2,\pm}\right|_{\mathrm{high}}$ $=-\mu\alpha^{\prime\prime}\pm\sqrt{\mu\epsilon+\mu^{2}\alpha^{\prime\prime 2}},$ (8)
@@
III Birefringence and Rotatory Power
The circular birefringence is measured in terms of the rotatory power,
$\delta_{RP}=-\dfrac{\omega}{2}\left[\operatorname{Re}(n_{LCP})\ -\operatorname{Re}(n_{RCP})\right],$ (9)
for waves propagating in the same direction. The fact that we have four refractive indices, associated with forward and backwards waves, allows us to write the following expressions for rotatory power:
@@ III Birefringence and Rotatory Power
$\delta_{-,-}$ $=\frac{\omega}{2}\operatorname{Re}\left(2\mu\alpha^{\prime\prime}+Q_{+}-Q_{-}\right),$ (11)
where the first subscript ($\pm$) stands for $n_{2,\pm}$, and the second one for $n_{1,\pm}$. Notice that one can not write $\delta_{-,+}$ since it is related to a backward RCP wave with $n_{2-}$ and a forward RCP wave with $n_{1,+}$. Furthermore, $\delta_{+,-}$ holds in the frequency range $\omega<\omega_{R}$, where the mode associated with $n_{1,-}$ corresponds to forward propagation of an RCP wave.
@@
The behavior of the rotatory power $\delta_{+,\pm}$ of Eq. (10) as a function of the frequency is observed in Fig. 3, being both equal in the range $0<\omega<\omega_{0}$ (see the purple line), where they only receive contribution [^1] One notes that $Q_{-}$ contributes to the RP only when it is real, that is, for $\omega>\omega_{0}$, with $\omega_{0}$ given by Eq. (6). of the first root ($Q_{+}$). In this frequency range, both RPs $\delta_{+,\pm}$ possess sign reversal at $\omega=\omega^{\prime}$, with $\omega^{\prime}$ given by $\omega^{\prime}={b}/({3\mu\alpha^{\prime\prime 2}-\epsilon})$. For $\omega>\omega_{0}$, the RPs begin to receive the contribution of the second root, resulting in the splitting starting at $\omega=\omega_{0}$, from which $\delta_{+,+}$ exhibits an increasing behavior (see the red line), while $\delta_{+,-}$ has decreasing magnitude (see the blue line). Along with that, the RP $\delta_{+,-}$ exhibits a second (additional) sign reversal at the frequency, $\omega^{\prime\prime}={b}/({2\alpha^{\prime\prime}\sqrt{\mu\epsilon}})$, with $\omega^{\prime\prime}>\omega_{0}$. This double sign inversion is an effect of both the magnetoelectric and AHE parameters, a very unusual feature that may work as an optical signature of this chiral system. A single RP reversion was reported for the case of a bi-isotropic dielectric in the presence of an isotropic axion magnetic current [43]. For high frequencies, where the term $b/\omega$ becomes negligible, the coefficients $\delta_{+,+}$ and $\delta_{+,-}$ recover a linear behavior with a positive (negative) slope.
@@
Figure 3: Rotatory power $\delta_{+,\pm}$ of Eq. (10), constructed using the two refractive indices for the RCP wave and $n_{2,+}$. The RP $\delta_{+,+}$ is depicted by the purple-red line, while $\delta_{+,-}$ is demarcated by the purple-blue line, endowed with double sign reversal. Here, we have used $\mu=1,\epsilon=\alpha^{\prime\prime}=3$, and $b=1$ $\mathrm{s}^{-1}$.
IV Kerr rotation and Kerr ellipticity
@@
Considering incident light stemming from medium 1 (a usual dielectric described by permittivity $\epsilon_{1}$ and permeability $\mu_{1}$) and reflecting on the surface of medium 2 [a chiral dielectric described by the parameters $\epsilon$, $\mu$, the axion vector $\mathbf{b}$, and the refractive indices of Eq. (4)], we obtain two Fresnel coefficients for each circularly polarized mode associated with $n_{1,\pm}$ and $n_{2,\pm}$ at normal incidence, namely
$r_{2,\pm}=\frac{\mu_{1}n_{2,\pm}-\mu n_{1}}{\mu_{1}n_{2,\pm}+\mu n_{1}},$ (12a)
$r_{1,\pm}=\frac{\mu_{1}n_{1,\pm}-\mu n_{1}}{\mu_{1}n_{1,\pm}+\mu n_{1}}.$ (12b)
The latter can be used to define the Kerr angle expressions [42]
$\Delta_{+,\pm}$ $=\left(\frac{r_{2,+}-r_{1,\pm}}{r_{2,+}+r_{1,\pm}}\right),$ (13)
$\Delta_{-,-}$ $=\left(\frac{r_{2,-}-r_{1,-}}{r_{2,-}+r_{1,-}}\right),$ (14)
@@
where the first (second) $\pm$ subscript is related to $n_{2,\pm}$ ($n_{1,\pm}$). The quantity $\Delta_{+,-}$ holds only in the frequency range $\omega<\omega_{R}$. The elements $\Delta_{+,\pm}$ and $\Delta_{-,-}$ are generally complex,
$\Delta_{+,\pm}=\Delta_{+,\pm}^{\prime}+i\Delta_{+,\pm}^{\prime\prime},$ (15)